# Towards Browser Controls to Protect Cookies from Malicious Extensions

### v2

@@ Appendix C Additional Discussion & Limitations
Reduced Memory Overhead: As mentioned in Section 7.1, both the Monitored changelog and report can grow rapidly for cookies with frequent changes. Thus one direction for future work is to minimize the size of these values. There are several ways this reduction may be achieved. In our design, both the Monitored log entries and message are C++ string objects. These objects add 24 bytes of overhead to each cookie and each changelog entry. Thus switching to another representation such as a basic character pointer could help reduce some of the storage overhead. Similarly, we already encode the changelog entries and current settings to minimize their size. However, as they are both strings it may be possible to use other string compression algorithms to reduce their size further. Finally, it may be possible to reduce which attributes are logged. Some attributes such as HttpOnly prevent cookies from being stolen, but a stolen Monitored cookie cannot be used outside the user’s browser. As such it may not be necessary to log changes made to these attributes.
@@
Further Extension Analysis: During our analysis of extensions, we specifically looked for which extensions could access cookies. However, during the process, we noted some other interesting results. During the initial scrape, we identified several nearly identical extensions. These extensions had identical names, highly similar descriptions and assets, and the majority of their code was the same. However, it is unclear why some many duplicates exist and how they differ. Similarly, while Chrome and Firefox both support the WebExtension APIs they do not always share the same implementation. Despite this many extensions work across browsers. Due to these observations, another interesting direction for future work is further analysis of these extensions to determine how they differ from each other, across browsers, and what effect this may have on the security of the browser.
@@ Appendix D Extension Security: Extended Related Work
Given their popularity and access to powerful APIs, several studies have analyzed extensions in terms of the APIs they commonly use, how they impact other controls, and how users understand their impacts [99, 98, 13, 30]. Additionally, prior work also investigates the privacy threats of extensions, such as tracking user behavior and interests [108]. As such many works investigate extension fingerprinting techniques [109, 110] and ways to prevent it [111, 112].
As mentioned in Section 1, extensions can be exploitable yielding extension permissions to $\sf{\mathcal{A}dv}$. Many static analysis techniques and browser modifications have been proposed to help detect/mitigate extension vulnerabilities [34, 113, 114, 33]. Several studies also acknowledge the threat posed by malicious extensions focusing on detection [25, 115, 24, 116, 117, 118, 26] whereas fewer prior attempts aim to prevent malicious extension actions. WebEnclave [35] is an extension that creates an isolated enclave on a website to hide secure elements such as login prompts from other extensions. Similarly, WRIT [36] uses service workers to attach tokens to JS events (i.e. mouse clicks) to verify if they were generated by the user rather than an extension. Picazo-Sanchez et. al. [37] create a browser extension to monitor changes made by each extension, remove them before the next extension runs, and re-add all the changes after all extensions have executed. The latter prevents a malicious extension from abusing execution orders to learn about/tamper with other extensions. These works focus on website interactions rather than broader browser APIs like Cookies and WebRequest.

### v3

@@ Appendix C Additional Discussion & Limitations
Reduced Memory Overhead: As mentioned in Section VII-A, both the changelog and Monitored report can grow rapidly for cookies with frequent changes. Thus one direction for future work is to minimize the size of these values. There are several ways this reduction may be achieved. In our design, both the Monitored log entries and messages are C++ string objects. These objects add 24 bytes of overhead to each cookie and each changelog entry. Thus switching to another representation such as a basic character pointer could help reduce some of this storage overhead. Similarly, we already encode the changelog entries and cookie’s current settings to minimize their size. However, as they are both strings it may be possible to use other string compression algorithms to reduce their size further. Finally, it may be possible to reduce which attributes are logged. Some attributes such as HttpOnly prevent cookies from being stolen, but a stolen Monitored cookie cannot be used outside the user’s browser. As such it may not be necessary to log changes made to these attributes.
@@
Further Extension Analysis: During our analysis of extensions, we specifically looked for which extensions could access cookies. However, during the process, we noted some other interesting patterns. During the scrape, we identified several nearly identical extensions. These extensions had identical names, highly similar descriptions and assets, and the majority of their code was the same. However, it is unclear why so many duplicates exist and how they differ. Similarly, while Chrome and Firefox both support the WebExtension APIs they do not always share the same implementation. Despite this many extensions work across browsers. Due to these observations, another interesting direction for future work is further analysis of these extensions to determine how they differ from each other, across browsers, and what effect this may have on the security of the browser.
@@ Appendix D Extension Security: Extended Related Work
Given their popularity and access to powerful APIs, several studies have analyzed extensions in terms of the APIs they commonly use, how they impact other controls, and how users understand their impacts [111, 112, 13, 34]. Additionally, prior work also investigates the privacy threats of extensions, such as tracking user behavior and interests [113]. As such many works investigate extension fingerprinting techniques [114, 115] and mitigations [116, 117].
As mentioned in Section I, extensions can be exploited to yield extension permissions to $\sf{\mathcal{A}dv}$. As such, many analysis techniques and browser modifications have been proposed to help detect/mitigate extension vulnerabilities [38, 118, 119, 37, 120]. Several studies also acknowledge the threat posed by malicious extensions focusing on detection [28, 121, 27, 122, 123, 124, 29] whereas fewer prior attempts aim to prevent malicious extension actions. WebEnclave [39] is an extension that creates an isolated enclave on a website to hide secure elements such as login prompts from other extensions. Similarly, WRIT [40] uses service workers to attach tokens to JS events (i.e. mouse clicks) to verify if they were generated by the user rather than an extension. Picazo-Sanchez et. al. [41] create a browser extension to monitor changes made by each extension, remove them before the next extension runs, and re-add all the changes after all extensions have executed. The latter prevents a malicious extension from abusing execution orders to learn about/tamper with other extensions. Unlike CREAM, these works focus on website interactions rather than broader browser APIs like Cookies and WebRequest.